# Convergence analysis of Levenberg-Marquardt method with Singular Scaling for non-zero residue nonlinear least-squares problems [^1] This work is supported by FAPESC (Fundação de Amparo à Pesquisa e Inovação do Estado de Santa Catarina) [grant number 2023TR000360]

### v1

@@ Abstract
This paper addresses the convergence analysis of a variant of the Levenberg-Marquardt method (LMM) designed for nonlinear least-squares problems with non-zero residue. This variant, called LMM with Singular Scaling (LMMSS), allows the LMM scaling matrix to be singular, encompassing a broader class of regularizers, which has proven useful in certain applications. In order to establish local convergence results, a careful choice of the LMM parameter is made based on the gradient linearization error, dictated by the nonlinearity and size of the residual. Under completeness and local error bound assumptions we prove that the distance from an iterate to the set of stationary points goes to zero superlinearly and that the iterative sequence converges. Furthermore, we also study a globalized version of the method obtained using line-search and prove that any limit point of the generated sequence is stationary. Some examples are provided to illustrate our theoretical results.
Keywords. Levenberg-Marquardt, Singular Scaling Matrix, Convergence analysis, Non-zero residue
@@ 1 Introduction
In this paper, we investigate global and local convergence properties of the Levenberg–Marquardt method applied to a nonlinear least-squares (NLS) problem:
$\min_{x\in\mathbb{R}^{n}}\quad\frac{1}{2}\|F(x)\|^{2}:=\phi(x),$ (1)
where $F:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is twice continuously differentiable. In particular, we focus on the overdetermined problem, where $m\geq n$.
Unlike previous research, which predominantly focused on the zero residue case or in cases where the rank of the Jacobian is complete [7, 6, 10, 11, 13], here we consider NLS problems where the set of zeros of $F$ is empty and its Jacobian is allowed to be rank-deficient.
We say that problem (1) is a non-zero-residue NLS problem when the set of zeros of $F$ is empty. Our interest is to identify the stationary points of $\phi$, i.e., the set $X^{*}=\{x\in\mathbb{R}^{n}\mid J(x)^{T}F(x)=0\}$, where $J(x)\in\mathbb{R}^{m\times n}$ is the Jacobian of $F$ at $x$, under the assumption that $X^{*}\neq\emptyset$. Throughout the text we shall use $x^{*}$ to denote an arbitrary element of $X^{*}$.
Although works such as [2, 3] explore problems with non-zero residuals and present results of local convergence of the LM method, it should be noted that the scale matrix used in these works is the identity matrix. In this work, in addition to assuming non-zero residue, we consider the following iteration:
@@
where $F_{k}:=F(x_{k})$, $J_{k}:=J(x_{k})$ is the Jacobian of $F$ at $x_{k}$, $\lambda_{k}>0$, $\alpha_{k}$ is the step size, and $L^{T}L$ referred to as scaling matrix, is allowed to be singular. We refer to the iteration (2)–(3) as the Levenberg-Marquardt method with Singular Scaling (LMMSS).
This type of iteration is motivated by applications in inverse problems, as recently seen in [5], where a parameter recovery problem using LMMSS is addressed, including a local convergence analysis for the zero residue case. On the other hand, convergence properties of the classic LMM in the non-zero residue scenario have received recent attention in the literature [2, 3]. These are two aspects in which our contribution differs from the existing literature.
Our aim is to analyze the local convergence of LMM with singular scaling matrices in the non-zero residue scenario. Furthermore, we shall provide a globalization for the LMMSS iteration where the step-size $\alpha_{k}$ is determined by line-search and prove that any limit point of the generated sequence is stationary for the least-squares function.
This paper is organized as follows. In Section 2, we start with some assumptions that will be considered throughout the manuscript. In Section 3, we present technical results necessary for the upcoming sections. We show that LMMSS iteration is well-defined and provide the local convergence results in Section 4, which are categorized into two cases based on the behavior of the rank of the Jacobian around a stationary point (constant and diminishing rank). Finally, Section 5 proves that, for LMMSS with an Armijo line-search scheme, every limit point of the generated sequence is a stationary point of the sum-of-squares function, thereby establishing the global convergence of this method. A few examples are given in Section 6 to illustrate our theoretical results. Final considerations are given in Section 7.
@@ 2 Assumptions
Recall that throughout this manuscript, we assume that the set of stationary points of $\phi$ is non-empty: $X^{*}\neq\emptyset$.
Assumption 1 below guarantees that $d_{k}$ solution of (2) is unique.
Assumption 1.
The matrix $L\in\mathbb{R}^{p\times n}$ has $\text{rank}(L)=p\leq n\leq m$ and there exists $\Omega\subset\mathbb{R}^{n}$ and $\gamma>0$ such that, for every $x\in\Omega$
@@
Such assumption comes from the inverse problems literature [12, 8] where it is sometimes called completeness condition. It was a key assumption in [5] to study the local convergence of LMMSS in the zero residue setting. We remark that such assumption is equivalent to $\mathcal{N}(J(x))\cap\mathcal{N}(L)=\{{\bf 0}\},\forall x\in\Omega.$
Next assumption is quite standard in the Gauss–Newton and Levenberg–Marquardt literature and asks the Jacobian to be locally Lipschitz.
Assumption 2.
For any $x^{*}\in X^{*}$, there exists a constant $\delta\in(0,1)$ and $L_{0}>0$ such that
$\|J(x)-J(y)\|\leq L_{0}\|x-y\|,$ (5)
@@
$\|J(y)(x-y)-(F(x)-F(y))\|\leq L_{1}\|x-y\|^{2},\quad\forall x,y\in B(x^{*},\delta)$ (6)
@@
Lemma 1.
[2, Lemma 2.1] If Assumption 2 is satisfied, then there exists a constant $L_{4}>0$ such that
@@
Next, we present the local error bound condition assumed in this work. We remark that it is an error bound condition with respect to the set of stationary points of $\phi$ rather than the zeroes of $F$ (which is empty in the case of non-zero residue).
Assumption 3 (Error bound).
For any $x^{*}\in X^{*}$, $\left\lVert J(x)^{T}F(x)\right\rVert$ provides an error bound in $B(x^{*},\delta)$, i.e., there exists $\omega\in(0,\infty)$ such that
$\omega\textup{dist}(x,X^{*})\leq\left\lVert J(x)^{T}F(x)\right\rVert,\quad\forall x\in B(x^{*},\delta),$ (10)
@@
Throughout the text, given $x$, we shall denote by $\bar{x}$ an element of $X^{*}$ such that $\|x-\bar{x}\|=\text{dist}(x,X^{*})$. Also, let $x^{*}\in X^{*}$ and $J_{*}=J(x^{*})$.
From Assumptions 3 and (7), we obtain
$\omega\text{dist}(x_{k},X^{*})\leq\left\lVert J_{k}^{T}F_{k}\right\rVert\leq L_{3}\text{dist}(x_{k},X^{*}).$ (11)
The remaining assumption focuses, as outlined in (8), on the terms
@@
These terms play a crucial role in controlling the error of the “incomplete linearization” of the gradient, as expressed in (9[2].
Assumption 4.
For every $x\in B(x^{*},\delta)$ and every $z\in X^{*}\cap B(x^{*},\delta)$, the following inequality holds:
$\|(J(x)-J(z))^{T}F(z)\|\leq C\|x-z\|^{1+r},$ (12)
with $r\in(0,1]$ and $C\geq 0$.
@@ 3 Auxiliary Results
In this section, we organize other fundamental results essential for the upcoming analysis.
We start by mentioning that (2) always has a solution. In fact, (2) corresponds to the normal equation $B^{T}Bd=B^{T}c$ of the augmented system
$Bd:=\left[\begin{array}[]{c}J_{k}\\ \sqrt{\lambda_{k}}L\end{array}\right]d=-\left[\begin{array}[]{c}F_{k}\\ 0\end{array}\right]=:c,$
and the normal equation always has a solution because $B^{T}c$ lies in the range of $B^{T}B$. But it is Assumption 1 that ensures the solution of (2) is unique. Indeed, it is easy to see that condition (4) at $x_{k}$ implies that $J_{k}^{T}J_{k}+\lambda_{k}L^{T}L$ is positive definite (and thus nonsingular).
Next result will be important when the rank of $J(x)^{T}J(x)$ is constant in a neighborhood of $x^{*}$.
Lemma 2.
[2, Corollary 3.1] Given $\kappa>1$, if
$\text{rank}(J(x)^{T}J(x))=\text{rank}(J_{*}^{T}J_{*})=q\geq 1,$
and
@@
Now, we will show some results on the Generalized Singular Value Decomposition (GSVD), an important tool in theoretical analysis of the LMMSS direction.
Theorem 1.
(GSVD) [9, p. 22] Consider the pair $(A,L)$, where $A\in\mathbb{R}^{m\times n}$, $L\in\mathbb{R}^{p\times n}$, $m\geq n\geq p$, $rank(L)=p$, and $\mathcal{N}(A)\cap\mathcal{N}(L)=\{0\}$. Then, there exist matrices $U\in\mathbb{R}^{m\times n}$ and $V\in\mathbb{R}^{p\times p}$ with orthonormal columns and a nonsingular matrix $X\in\mathbb{R}^{n\times n}$ such that
$A=U\left[\begin{array}[]{cc}\Sigma&0\\ 0&I_{n-p}\end{array}\right]X^{-1}\quad\text{and}\quad L=V\left[\begin{array}[]{cc}M&0\end{array}\right]X^{-1},$ (14)
@@
$\Sigma=\text{diag}(\sigma_{1},\dots,\sigma_{p})\in\mathbb{R}^{p\times p}\quad\text{and}\quad M=\text{diag}(\mu_{1},\dots,\mu_{p})\in\mathbb{R}^{p\times p}.$ (15)
Moreover, the elements of $\Sigma$ and $M$ are non-negative, ordered as follows:
$0\leq\sigma_{1}\leq\dots\leq\sigma_{p}\leq 1\quad\text{and}\quad 1\geq\mu_{1}\geq\dots\geq\mu_{p}>0,$ (16)
@@
Next result from [5] gives an upper bound on $\|X_{k}\|$ depending on $\gamma$ from Assumption 1.
Lemma 3.
[5, Lemma 2.2] Consider $\{x_{k}\}$ the sequence generated by LMMSS along with the GSVD of the pair $(J_{k},L)$. Then, $\forall k\in\mathbb{N}$, $\|X_{k}\|\leq 1/\sqrt{\gamma}$.
Remark 2.
@@
    For $\lambda\in(0,1)$, the function $\psi(\gamma,\lambda)$ has a unique maximum value attained at
@@
    For a fixed $\lambda\in[1,+\infty)$, the function $\psi(\gamma,\lambda)$ is non-decreasing and upper-bounded. More precisely, $\psi(\gamma,\lambda)\leq 1,\quad\forall\gamma\geq 0.$
Proof.
@@
$\lim_{\gamma\rightarrow\infty}\psi(\gamma,\lambda)=1.$ (21)
@@
    Consider $\lambda\in(0,1)$ fixed. As the maximizers of $\psi(\gamma,\lambda)$, for $\gamma>0$, must satisfy $\frac{\partial\psi}{\partial\gamma}(\gamma,\lambda)=0,$ from (22), we have $02(\lambda-1)\gamma=$ 0, thus $\gamma_{\max}=0$. Substituting $\gamma_{\max}$ into $\psi(\gamma,\lambda)$, we obtain
    $\psi(\gamma_{\max},\lambda)=1/2.$
@@
    Now, for fixed $\lambda\in[1,+\infty)$, from (22), we have $\frac{\partial\psi}{\partial\gamma}(\gamma,\lambda)>0$ for $\gamma\geq 0$. Therefore, $\psi$ is increasing for $\gamma\in[0,+\infty)$, with $\psi(0,\lambda)=0$. Then, from (21), we conclude that $\psi(\gamma,\lambda)\leq 1$, for all $\gamma\geq 0$.
We are now ready to prove two key lemmas for the local convergence analysis. They show that the norm of LMMSS direction $d_{k}$ is ${\cal O}(\textup{dist}(x_{k},X^{*}))$, considering two distinct cases. Lemma 5 addresses the scenario where the rank of the Jacobian around the solution (stationary-point) is constant, while Lemma 6 deals with the case of diminishing rank.
Lemma 5.
Suppose that Assumptions 1-3 are valid in $B(x^{*},\delta)$, and
@@
$\|d_{k}\|\leq c_{1}\textup{dist}(x_{k},X^{*}).$ (23)
Proof.
@@
which implies that $\left\lVert J_{k}d_{k}\right\rVert^{2}+\lambda_{k}\left\lVert Ld_{k}\right\rVert^{2}=-(J_{k}d_{k})^{T}F_{k}$, thus, if we assume that $d_{k}\in\mathcal{N}(J_{k})$, we conclude that $\lambda_{k}\left\lVert Ld_{k}\right\rVert^{2}=0$. But $\left\lVert Ld_{k}\right\rVert>0$, given that, by Assumption 1, $\mathcal{N}(J_{k})\cap\mathcal{N}(L)=0$, and also $\lambda_{k}\neq 0$, and we have a contradiction.
@@
Let $J_{k}=U\Sigma V^{T}$ be the SVD for $J_{k}$. As $d_{R}\in\mathcal{N}(J_{k})^{\perp}$, $d_{R}$ can be written as a linear combination of the columns of $V$ corresponding to the non-zero singular values of $J_{k}$. Therefore,

### v2

# Florianópolis, 88040900, SC, Brazil

@@ Abstract
Recently, a Levenberg-Marquardt method with Singular Scaling matrix, called LMMSS, was proposed and successfully applied in parameter estimation in heat conduction problems, where the choice of suitable singular scaling matrix resulted in better quality approximate solutions than those of the classical Levenberg-Marquardt. In this paper, we study convergence properties of LMMSS when applied to nonzero residual nonlinear least-squares problems. We show that the local convergence of the iterates depends both on the control of the gradient linearization error and on a suitable choice of the regularization parameter. Incidentally, we show that the rate of convergence is dictated by a measure of nonlinearity and residual size, so that if such a measure goes to zero quickly enough, the convergence can be superlinear, otherwise, in general, we show that not even linear convergence can be expected if such a measure is not small enough. Additionally, we propose a globalized version of the method and prove that any limit point of the generated sequence is stationary for the least-squares function. Some examples are provided to illustrate our theoretical results.
Keywords:
Levenberg-MarquardtSingular Scaling Matrix Convergence analysis nonzero residual
MSC:
49M37 65K05 90C30
@@ 1 Introduction
In this study we investigate convergence properties of the Levenberg–Marquardt method with singular scaling (LMMSS) from Boos2024 when applied to the nonlinear least-squares (NLS) problem:
$\min_{x\in\mathbb{R}^{n}}\quad\phi(x):=\frac{1}{2}\|F(x)\|^{2},$ (1)
[…]
Unlike previous local convergence analyses Dennis_Schnabel ; FanYuan ; Gratton2007 ; Yamashita , including Boos2024 , we assume neither zero residual at a solution of (1) nor full rank of the Jacobian at such a point. In applied contexts, such as data fitting, parameter estimation, experimental design, and imaging problems bellavia2018elliptical ; barz2015nonlinear ; cornelio2011regularized ; deidda2014regularized ; henn2003levenberg ; landi2017limited , to name a few, admitting a nonzero residual is essential for achieving meaningful solutions.
The LMMSS iteration is defined as follows Boos2024 :
@@
where $F_{k}:=F(x_{k})\in\mathbb{R}^{m}$, $J_{k}:=J(x_{k})\in\mathbb{R}^{m\times n}$ is the Jacobian of $F$ at $x_{k}$, $\lambda_{k}>0$, $\alpha_{k}$ is the step size and $L^{T}L$, referred to as scaling matrix, is allowed to be singular. We refer to the iteration (2)–(3) as the Levenberg-Marquardt method with Singular Scaling (LMMSS). When $L^{T}L=I$ we retrieve the classic Levenberg-Marquardt method (LMM).
In general, problem (1) is a nonconvex optimization problem and we will limit our attention to stationary points of $\phi$. The set of stationary points will be denoted by
$X^{*}=\{x\in\mathbb{R}^{n}\mid\nabla\phi(x)=J(x)^{T}F(x)=0\},$
and assume that $X^{*}\neq\emptyset$. We are particularly interested in the case of nonisolated stationary points, with a possible change (decrease) in the rank of the Jacobian as the generated sequence of iterates $\{x_{k}\}$ approaches the set $X^{*}$.
Given a starting point sufficiently close to $x^{*}\in X^{*}$, we are interested in the convergence analysis of the sequence generated by the local LMMSS iteration (with $\alpha_{k}=1$). Our main contribution consists in establishing local convergence results for LMMSS based on an error bound condition upon $\nabla\phi$ without requiring zero residual neither full rank of the Jacobian at stationary points.
[…]
The seminal work Yamashita showed the local convergence of LMM for systems of nonlinear equations under an error bound condition upon the norm of the residual, without relying on the assumption of the full rank of the Jacobian. Assuming $\lambda_{k}=\|F(x_{k})\|^{2}$, the authors established that the distance of the iterates to the solution set converges quadratically to zero. Later, FanYuan improved such result and showed that quadratic convergence is still attained for $\lambda_{k}=\|F(x_{k})\|^{1+r},r\in[0,1]$. Furthermore, they also showed that the sequence of iterates itself converges quadratically to some point in the solution set.
More recently, a few papers have focused on establishing local convergence of LMM for nonzero residual NLS problems (i.e $\forall x\in\mathbb{R}^{n}$, $F(x)\neq 0$). In behling2019local , local convergence was established using an error bound on the gradient of $\phi$ and considering a possible change in the rank of Jacobian around a stationary point. In Bergou2020 , local convergence of LMM in the nonzero residual case was also established, but using a different error bound condition based on the distance between the residual vector at $x$ and the one at $\bar{x}$, the stationary point closest to $x$.
Nevertheless, all the works mentioned above considered the classic LMM where $L^{T}L=I$. Perhaps Boos2024 was the first work to consider local convergence analysis of LMMSS (for $L^{T}L\neq I$) under an error bound condition, but considering the zero residual case. The motivation to use singular matrices in the form $L^{T}L$ in LMMSS iteration (2)–(3) comes from the general form of Tikhonov regularization Tikhonov95 for linear systems $Ax=b$, i.e.,
$x_{\lambda}=(A^{T}A+\lambda L^{T}L)^{-1}A^{T}b.$
The use of a singular scaling matrices allows us to use seminorm regularizers which can promote specific properties (inherent of the expected solution) in the approximate solution generated by method. Like the $\ell_{1}$-norm is used to promote sparsity, if the aimed solution is expected to be smooth, then the use of $L$ as a discrete version of derivative operator allows us to induce such smoothness in the approximate solution. This approach proved to be successful in ill-posed linear inverse problems Hansen1998 ; Hansen2010 ; Buccini2017 .
Nonlinear inverse problems can also benefit from the more flexible choice of $L$ in LMMSS. For instance, in Boos2024 LMMSS was successfully used to estimate the 2D perfusion coefficient in a bioheat model. Figure 1 shows the surface plots of the 2D perfusion coefficient reconstructions: the picture on the left plots the expected solution, the one on the middle shows the reconstruction obtained by the classic LMM and the one on the right corresponds to the reconstruction obtained by LMMSS with a specific choice of $L=\mathcal{L}_{3}$, where $\mathcal{L}_{3}$ is based on a discrete version of the third order derivative operator. Both methods stopped after dozens of iterations based on the discrepancy principle as stopping criterion Morozov1984 . For this problem, it is noteworthy that although the temperature reconstruction error was the same for both methods (almost the same value of $\phi$), LMMSS provided an approximate solution closer to the expected one. For more details, see (Boos2024, , Section 5.1)).
Figure 1: The picture on the left plots the expected solution, the one on the middle shows the reconstruction obtained by the classic LMM and the one on the right corresponds to the reconstruction obtained by LMMSS with a specific choice of $L$. Image extracted from (Boos2024, , Figure 3).
The above example illustrates the potential of LMMSS when a suitable $L$ is chosen. Observe that $L$ defines the seminorm regularizer and must be chosen according to the problem at hand, based on known properties of the expected solution. As we already mentioned, if the solution (of the original problem, whose discretization gave rise to the nonlinear least-squares problem) is expected to be smooth, then the use of $L$ as a discrete version of derivative operators may be a good choice. But other examples are available in the literature, for instance, see Buccini2017 ; pes2020 ; Pes2022 and references therein.
As far as we know, convergence analysis of LMMSS in the nonzero residual scenario has not yet been addressed in the literature. This is precisely the purpose of this work: for the LMMSS iteration defined in Boos2024 we aim at establishing local convergence rates to stationary points of the least-squares function in the case of nonzero residual under an error bound condition upon $\nabla\phi$, following the lines of the analysis in behling2019local . Of course, by considering a possibly singular $L^{T}L$ in iteration (2)–(3) makes the analysis more challenging, for the iteration itself to be well defined depends on a condition on the null spaces of $L$ and the Jacobian of $F$.
Additionally, we propose a globalization for the LMMSS iteration where the step-size $\alpha_{k}$ is determined by a line-search satisfying an Armijo condition and the search direction is safeguarded by the classic LMM direction. For the resulting algorithm, we prove that any limit point of the generated sequence is stationary.
The paper is organized as follows. In Section 2, we gather some necessary assumptions to show that LMMSS iteration is well-defined and for the subsequent analysis. In Section 3, we present some technical results and discuss the local convergence analysis in Section 4. As we shall see, the convergence rate is dictated by a measure of nonlinearity and residual size. If such a measure goes to zero fast enough, the convergence can be superlinear but, in general, we cannot expect even linear convergence if such measure is not small enough. In Section 5, a globalized LMMSS algorithm is proposed and its global convergence established. To close this theoretical study, some examples are given in Section 6 to illustrate the local and global behavior of the algorithm in nonzero residual nonlinear least-squares problems. Final considerations are drawn in Section 7.
@@ 2 Assumptions
Recall that throughout this manuscript, we assume that $X^{*}\neq\emptyset$. We shall denote a generic element of $X^{*}$ by $x^{*}$. Given $x$, we denote by $\bar{x}$ an element of $X^{*}$ such that $\|x-\bar{x}\|=\text{dist}(x,X^{*})$. Also, given $x^{*}\in X^{*}$, we may use the shorthand notation $F_{*}=F(x^{*})$ and $J_{*}=J(x^{*})$. The nullspace of $A$ is denoted by $\mathcal{N}(A)$ and $\mathcal{R}(A)$ denotes its range.
Definition 1
We say that the uniform completeness condition for problem (1) with respect to $L$ holds at $\Omega\subset\mathbb{R}^{n}$, if $L\in\mathbb{R}^{p\times n}$ has $\text{rank}(L)=p\leq n\leq m$ and there exist a constant $\gamma>0$ such that, for every $x\in\Omega$
@@
It is worth noting that condition (4) is equivalent to
$\mathcal{N}(J(x))\cap\mathcal{N}(L)=\{{\bf 0}\}.$ (5)
Condition (5) is often referred to as a completeness condition and comes from the literature of linear inverse problems Morozov1984 ; engl , where $J(x)=A,\forall x\in\mathbb{R}^{n}$. Observe that it clearly holds when $J(x)$ is of full rank, regardless of the choice of $L$, as it occurs in some applications involving physical problems (Boos2024, , Section 5.1).
In the context of LMMSS iteration, if the iterates $x_{k}$ lie in a set $\Omega\subset\mathbb{R}^{n}$ where the uniform completeness condition holds, then $d_{k}$ solution of (2) exists and is unique because (4) implies $J_{k}^{T}J_{k}+\lambda_{k}L^{T}L$ is positive definite for $\lambda_{k}>0$.
Next assumption is standard in the Gauss–Newton and Levenberg–Marquardt literature and asks the Jacobian to be locally Lipschitz.
Assumption 1
For some $x^{*}\in X^{*}$, there exists a constant $\delta\in]0,1[$ and $L_{0}>0$ such that
$\|J(x)-J(y)\|\leq L_{0}\|x-y\|,$
@@
$\|J(y)(x-y)-(F(x)-F(y))\|\leq L_{1}\|x-y\|^{2},\quad\forall x,y\in B(x^{*},\delta)$
@@
Lemma 1
(behling2019local, , Lemma 2.1) If Assumption 1 is satisfied, then there exists $\delta\in]0,1[$ and a constant $L_{4}>0$ such that
@@
Next, we present a local error bound condition assumed in this work.
Assumption 2
For some $x^{*}\in X^{*}$, $\left\lVert J(x)^{T}F(x)\right\rVert$ provides a local error bound at $x^{*}$, i.e., there exists $\delta\in]0,1[$ and $\omega>0$ such that
$\omega\textup{dist}(x,X^{*})\leq\left\lVert J(x)^{T}F(x)\right\rVert,\quad\forall x\in B(x^{*},\delta),$
@@
Error bound conditions have been extensively studied and used in convergence analysis of LM methods in the last two decades Yamashita ; fan ; FanYuan ; Fischer2002 ; behling2012 ; behling2013 ; behling2019local ; Fischer2024 . For zero residual problems, it is common to use $\|F(x)\|$ as an error bound. However, for nonzero residual problems, $F(x)\neq 0,\forall x\in\mathbb{R}^{n}$, and Assumption 2 uses $\|\nabla\phi(x)\|$ as an error bound instead. Section 6 presents some examples where such assumption holds. See also the discussion and examples in behling2019local . It is well-known that Assumption 2 is suitable to deal with nonisolated stationary points and it is weaker than the assumption of the Jacobian having full rank at $x^{*}\in X^{*}$. By the way, $J(x^{*})$ having full rank implies Assumption 2.
From Assumption 2 and (6), we obtain
$\omega\text{dist}(x,X^{*})\leq\left\lVert J(x)^{T}F(x)\right\rVert\leq L_{3}\text{dist}(x,X^{*}).$ (9)
The remaining assumptions focus, as outlined in (7), on the terms
@@
These terms play a crucial role in controlling the error, as expressed in (8), of the “incomplete linearization” of the gradient: notice that $J(x)^{T}J(x)$ is used in (8) instead of the Hessian $\nabla^{2}\phi(x)=J(x)^{T}J(x)+S(x)$, with $S(x)=\sum_{i=1}^{m}F_{i}(x)\nabla^{2}F_{i}(x)$, that would appear in a first order Taylor approximation of $\nabla\phi$.
Assumption 3
For some $x^{*}\in X^{*}$, for all $x\in B(x^{*},\delta)$ and all $z\in X^{*}\cap B(x^{*},\delta)$, the following inequality holds:
$\|J(x)^{T}F(z)\|\leq\sigma\|x-z\|,$ (10)
with $0\leq\sigma<\bar{\sigma}$, where $\bar{\sigma}$ is a positive constant depending on $\omega$, $L_{3}$ and the smallest positive eigenvalue of $J(x^{*})^{T}J(x^{*})$.
Explicit expressions for $\bar{\sigma}$ will be given ahead. Assumption 3 dates back, at least, to the work of Dennis Dennis1977 , and it was also considered in behling2019local where the authors analyze the local convergence of the standard LM (where $L^{T}L=I$) in the case of nonzero residual. It is a condition to control the linearization error of the gradient. In Dennis1977 , convergence analysis of the classic LMM is given for the nonzero residual case, under the assumption that $J(x^{*})^{T}J(x^{*})$ is nonsingular and $\|J(x)^{T}F(x^{*})\|\leq\sigma\|x-x^{*}\|$, in a neighborhood of $x^{*}\in X^{*}$, for a sufficiently small $\sigma>0$, namely $\sigma<\lambda^{*}:=\lambda_{\min}(J(x^{*})^{T}J(x^{*}))$. Here, because we are not requiring $J(x^{*})$ to have full column rank, the set $X^{*}\cap B(x^{*},\delta)$ may be a nonisolated set of stationary points and for this reason Assumption 3 is slightly different from that of Dennis1977 .
Remark 1
Observe that
$\|(J(x)-J(z))^{T}F(z)\|\leq\|J(x)-J(z)\|\|F(z)\|\leq L_{0}\|F(z)\|\|x-z\|.$
Thus, given a bound $\|F(z)\|\leq\beta$ for every $z\in X^{*}\cap B(x^{*},\delta)$, if $L_{0}\beta<\bar{\sigma}$, then (10) holds with $\sigma=L_{0}\beta$. This might occur when the residual is small enough in $X^{*}\cap B(x^{*},\delta)$.
Assumption 4
For some $x^{*}\in X^{*}$, for all $x\in B(x^{*},\delta)$ and all $z\in X^{*}\cap B(x^{*},\delta)$, the following inequality holds:
$\|J(x)^{T}F(z)\|\leq C\|x-z\|^{1+r},$
with $r\in]0,1]$ and $C\geq 0$.
Assumption 4 is admittedly stronger than Assumption 3, but it might hold in some specific problems. For example, it is clear that such assumption holds when $F$ is linear or $F(z)=0$. In behling2019local the authors showed that when Assumption 4 holds, LMM local convergence rate can be superlinear under a suitable choice of the LM parameter. In Dennis_Schnabel , it is mentioned that $J(x)^{T}F(z)\approx S(z)(x-z)$ and thus $\sigma$ could be interpreted as a combined measure of nonlinearity and residual size. Observe that $S(z)=0$ is also a sufficient (but not necessary) condition to ensure Assumption 4. See the examples in Section 6 and in the reference behling2019local .
@@ 3 Auxiliary Results
In this section, we organize other fundamental results for the upcoming analysis.
We start with a perturbation lemma that will be important when the rank of $J(x)^{T}J(x)$ is constant in a neighborhood of $x^{*}$.
Lemma 2
(behling2019local, , Corollary 3.1) Given $\kappa>1$, if $\text{rank}(J(x)^{T}J(x))=\text{rank}(J_{*}^{T}J_{*})=q\geq 1$, and
@@
Now, we will recall some results on the Generalized Singular Value Decomposition (GSVD), an important tool in theoretical analysis of the LMMSS direction.
Theorem 1
(GSVD) (Hansen1998, , p. 22) Consider the pair $(A,L)$, where $A\in\mathbb{R}^{m\times n}$, $L\in\mathbb{R}^{p\times n}$, $m\geq n\geq p$, $rank(L)=p$, and $\mathcal{N}(A)\cap\mathcal{N}(L)=\{0\}$. Then, there exist matrices $U\in\mathbb{R}^{m\times n}$ and $V\in\mathbb{R}^{p\times p}$ with orthonormal columns and a nonsingular matrix $X\in\mathbb{R}^{n\times n}$ such that
$A=U\left[\begin{array}[]{cc}\Sigma&0\\ 0&I_{n-p}\end{array}\right]X^{-1}\quad\text{and}\quad L=V\left[\begin{array}[]{cc}M&0\end{array}\right]X^{-1},$
@@
$\Sigma=\text{diag}(\sigma_{1},\dots,\sigma_{p})\in\mathbb{R}^{p\times p}\quad\text{and}\quad M=\text{diag}(\mu_{1},\dots,\mu_{p})\in\mathbb{R}^{p\times p}.$
Moreover, the elements of $\Sigma$ and $M$ are nonnegative, ordered as follows:
$0\leq\sigma_{1}\leq\dots\leq\sigma_{p}\leq 1\quad\text{and}\quad 1\geq\mu_{1}\geq\dots\geq\mu_{p}>0,$
@@
The following result provides bounds for the matrix $X_{k}$.
Lemma 3
(Boos2024, , Lemma 2.2) Consider $\{x_{k}\}$ the sequence generated by LMMSS along with the GSVD of the pair $(J_{k},L)$. If $\{x_{k}\}\subset\Omega$ where the uniform completeness condition holds, then for each $k\in\mathbb{N}$, we have
$\|X_{k}\|\leq\dfrac{1}{\sqrt{\gamma}}.$ (15)
Remark 3
@@
    For $\lambda\in]0,1[$, the function $\psi(\gamma,\lambda)$ has a unique maximum value attained at
@@
    For a fixed $\lambda\in[1,+\infty)$, the function $\psi(\gamma,\lambda)$ is nondecreasing and upper-bounded. More precisely,
    $\psi(\gamma,\lambda)\leq 1,\quad\forall\gamma\geq 0.$
Proof
@@
$\lim_{\gamma\rightarrow+\infty}\psi(\gamma,\lambda)=1=\lim_{\gamma\rightarrow-\infty}\psi(\gamma,\lambda).$ (17)
@@
    Consider $\lambda\in]0,1[$ fixed. As the maximizers of $\psi(\gamma,\lambda)$, for $\gamma\in\mathbb{R}$, must satisfy $\frac{\partial\psi}{\partial\gamma}(\gamma,\lambda)=0,$ from (18), we have
    $2(\lambda-1)\gamma=0\quad\Rightarrow\quad\gamma_{\max}=0,$
    Substituting $\gamma_{\max}$ into $\psi(\gamma,\lambda)$, we obtain
    $\psi(\gamma_{\max},\lambda)=\frac{1}{\lambda}.$
@@
    Now, for fixed $\lambda\in[1,+\infty)$, from (18), we have $\frac{\partial\psi}{\partial\gamma}(\gamma,\lambda)>0$ for $\gamma>0$. Therefore, $\psi$ is increasing for $\gamma\in[0,+\infty)$, with $\psi(0,\lambda)=1/\lambda\leq 1$. Then, from (17), we conclude that $\psi(\gamma,\lambda)\leq 1$, for all $\gamma\geq 0$. ∎
We are now ready to prove two key lemmas for the local convergence analysis. They show that the norm of LMMSS direction $d_{k}$ is $O(\textup{dist}(x_{k},X^{*}))$, considering two distinct cases. Lemma 5 addresses the scenario where the rank of the Jacobian around a stationary-point is constant, while Lemma 6 deals with the case when the rank decreases [^2] Observe that the rank function is lower semicontinuous thus, when a sequence $x_{k}$ in a neighborhood of $x^{*}\in X^{*}$ approaches $x^{*}$, the rank of the Jacobian $J(x_{k})$ either remains constant or decreases, these are the only two possibilities..
Lemma 5
Suppose that Assumption 1 is valid in $B(x^{*},\delta)$, that the uniform completeness condition holds at $B(x^{*},\delta)\setminus X^{*}$, and
@@
$\|d_{k}\|\leq c_{1}\textup{dist}(x_{k},X^{*}).$
Proof
@@
which implies that $\left\lVert J_{k}d_{k}\right\rVert^{2}+\lambda_{k}\left\lVert Ld_{k}\right\rVert^{2}=-(J_{k}d_{k})^{T}F_{k}$, thus, if we assume that $d_{k}\in\mathcal{N}(J_{k})$, we conclude that $\lambda_{k}\left\lVert Ld_{k}\right\rVert^{2}=0$. But $\left\lVert Ld_{k}\right\rVert>0$, given that, by condition 4, $\mathcal{N}(J_{k})\cap\mathcal{N}(L)=0$, and also $\lambda_{k}\neq 0$, and we have a contradiction.
@@
Let $J_{k}=U\Sigma V^{T}$ be the SVD for $J_{k}$. As $d_{R}\in\mathcal{N}(J_{k})^{\perp}$, $d_{R}$ can be written as a linear combination of the columns of $V$ corresponding to the nonzero singular values of $J_{k}$. Therefore,